

********************************************************************
UARS WEEKLY STATUS REPORT - 19 MAR 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully performed a major calibration using all four
deuterium lamps from March 12-15. As a part of this calibration
SUSIM's lower spectrometer, for the first time, scaned the sun to
provide additional and redundant calibration. At the same time, the
elements in the working optical channel were replaced to provide more
responsivity which should improve the noise characteristics of the
solar UV irradiance data.
With one exception, SUSIM gathered its normal solar, offset, and
occultation data without apparent problems. On March 13, the primary
long wavelength electrometer detector (PLE) entered a temporary
condition which prevented its proper calibration and normal operation.
This resulted in two lost scans. This abnormal behavior has not
recurred during subsequent usage. No operations errors have been
found that would explain this anomalous behavior. The cause remains
under investigation.
Reprocessing of the SUSIM data to data version 20 continues.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating limits. Science data products are of excellent
quality.
***********
MLS
***********
The instrument is now off as scheduled. MLS is scheduled to begin north viewing
operations on Apr 4 and operate though Apr 10.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII reprocessing is under way, with Version V5_11 of the
processing software.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
2 days CLA (7.0) ENG (L1-L3) (4R-2C-2F=2L)
49 days PEM (3.32) PROD (OA-L2) (760R-278C-1F=481L)
138 days SUS (7.0) PROD (668R-539C=129L)
131 days WIN (5.11) PROD (LOG) (2014R-131C=1883L)
13 days WIN (5.11) PROD (L1-L3) (2014R-13C=2001L)
The CDHF is scheduled to participate in an End to End Y2K test on
April 5th or April 19. This test will require a system clock time shift
and an interruption to normal operations.
Solstice excessive wallclock issue. No activity due to facilitiy
priority activities. Target next week to run in debug!
The CDHF has promoted WINDII V5.11 to production mode and started
mission reprocessing.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
This report covers 05 March 1999 (Orbit 40885, GMT Day 064) through 12 March
1999.
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 065/0116 40887 Full Time
HALOE 066/1212 40908 Off
HALOE 066/1351 40909 On
HALOE 069/0447 40948 Off
HALOE 069/0627 40949 On
HRDI 071/0006 40960 Day Only
PEM 071/0314 40977 AXIS On
The beta angle ranged between the angles of 46.2 degrees to 51.3 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 0 degrees on 28 March 1999 (DOY 087).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 05 March 1999 and 12 March 1999 are provided in
Appendix D.
The clock error ranged from -10.8 to 11.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% HALOE 04
% PEM 01
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/054 - 2287.496948
Transponder B: 99/055 - 2287.498731
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/070 show the observatory data loss to be
22 hours, 46 minutes, 32.756 seconds (an increase of 00.000 seconds since
last report period). This is a 0.0348 percent data loss which equals a
99.9652 percent data capture for the mission.
The UARS Test and Training Simulator (UTTS) remains non-operational. The
estimated time to repair the UTTS and return it to operation is unknown.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error. PEM EP6 is still not coming up nominally under
PMON control. The FOT will conduct additional tests to identify where in
the turn-on sequence the conflict occurs in the coming week.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
This task has been re-assigned as a part of the CSOC contract. FOT support
is currently suspended.
MEETINGS
UARS supported the first Monthly Project Meeting held on 10 March 1999.
Representatives from Central Data Handling Facility (CDHF), Flight Dynamics
Facility (FDF), Flight Operations Team (FOT), Y2K testing, CSOC customer
services and NASA GSFC (Project Scientist and Mission Director) were
present.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 28
March 1999 (DOY 087).
The spacecraft Roll attitude offset will be removed on DOY 075.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 March 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
40893 065/1157 99-092 OBC Clk Adj fr 20.6 to 0.00 msec/day
40908 066/1212 99-094 HALOE Power OFF
40909 066/1351 99-095 HALOE Power ON
40918 067/0453 99-096 OBC Clk Adj fr 0.00 to 20.6 msec/day
40948 069/0447 99-097 HALOE Power OFF
40949 069/0627 99-098 HALOE Power ON
40965 070/0812 99-099 OBC Clk Adj fr 20.6 to 0.00 msec/day
40970 070/1551 99-100 SUSIM LOW GRATING
40977 071/0314 99-101 PEM "Power ON, AXIS "
40988 071/2058 99-102 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 March 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
NBRECBI Y-HI NBTR B Current 40922,37,87
HATBAT Y-HI HALOE TBAT TMP 40939,78,80
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 March 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
40893 065/12:03 (06 Mar 99) from 20.6 to 00.0 msec/day
40918 067/04:57 (07 Mar 99) from 00.0 to 20.6 msec/day
40965 070/08:15 (11 Mar 99) from 20.6 to 00.0 msec/day
40988 071/21:07 (12 Mar 99) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 March 1999
05 Mar 1999 - GMT Day 064 Beta = 46.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -11.2/-33.6 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-16.8 +5.15/+3.87 +28.0/-15.2 26.2 1.025 26.0
3 +39.2/-11.2 +3.87/+2.92 +27.2/-16.0 26.1 1.021 25.7
12 Mar 1999 - GMT Day 071 Beta = 51.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -22.4/-44.8 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +44.8/-5.6 +5.47/+4.51 +26.8/-15.6 25.0 1.022 20.8
3 +56.0/-5.6 +4.51/+3.23 +26.4/-16.0 25.0 1.022 20.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 March 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-011 FOT 40887 065/0306 14170 Generic Late Acquisition
99-011 FOT 40934 068/0624 14170 Generic Late Acquisition
99-011 FOT 40987 071/1846 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-006 FOT see list Feb 99 14170 Generic Late Acq
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-011 FOT see list Mar 99 14170 Generic Late Acq
SEVNTFW99/11